OpenStudy (anonymous):

Jamie made a wax model of a rolling pin of diameter 8 cm. The rolling pin was shaped like a right circular cylinder with a right circular cone at each end as shown below. What was the total surface area of the rolling pin? Using complete sentences, describe the steps you used to calculate the surface area.

OpenStudy (anonymous):

wlcome to OS @liliana13lilly :

OpenStudy (anonymous):

thank you can you help me @bhaskar.pathk

OpenStudy (anonymous):

PLEASE!! lol

OpenStudy (anonymous):

surface area of cyllinder will be pi r^2 H where h is the hight of cyllinder ie 12 cm in your case so surface area of Cyllindrical part will be 3.14 x 81 x 12

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

surface area of cone is 1/3\[\pi \times r^2 \times h\] where according to your question h will be \[\sqrt{5^2 - 4^2 }\]

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

sorry in 1st case radious will be 4 cm so suface area will be 3.14 x 16 x 12 --- \[\pi \times r^2 \times h\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

given that diameter is 8 cm do you got how i taken the height of cone damm intrnet i too slow for me today...so it may be confusing for you

OpenStudy (anonymous):

no

OpenStudy (anonymous):

and its okay

OpenStudy (anonymous):

enjy with Os as it aim to solve your problem Rather than Givin you answer..